OpenStudy (anonymous):

If you double a number then add 24 you get 2/7 of the original number. What is the original number?

OpenStudy (anonymous):

set it up as an equation then solve

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

equation is \[2x+24=\frac{2}{7}x\]

OpenStudy (anonymous):

solve via \[2x-\frac{2}{7}x=-24\] \[\frac{12}{7}x=-24\] \[x=-24\times \frac{7}{12}\] \[x=-14\]

OpenStudy (anonymous):

what aq said